# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v1

@@ 2.2 Pontryagin’s Maximum Principle with a State-Dependent Market and Costate for $\mathbf{Y}_{t}$
Here, $\mathbf{Z}_{t}^{\lambda X}$ is represented as a $1\times n$ row vector.
Using Ito’s lemma to relate $\mathbf{Z}_{t}^{\lambda X}$ to the value function derivatives ($\lambda_{t}=V_{X}(t,X_{t},\mathbf{Y}_{t})$, so $\partial_{x}\lambda_{t}=V_{XX}$, $\partial_{\mathbf{Y}}\lambda_{t}=V_{X\mathbf{Y}}$), one can solve the FOC (4) for the optimal portfolio $\bm{\pi}_{t}^{*}$. The resulting expression, consistent with standard results in continuous-time finance, is:
@@
$\begin{split}\bm{\pi}_{t}^{*}=-\frac{\lambda_{t}^{*}}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}&\Sigma(t,\mathbf{Y}_{t}^{*})^{-1}\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t}^{*})-r(t,\mathbf{Y}_{t}^{*})\mathbf{1}\bigr{)}\\ -\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})}&\Sigma(t,\mathbf{Y}_{t}^{*})^{-1}\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\rho\bm{\sigma}_{Y}(t,\mathbf{Y}_{t}^{*})^{\top}(\partial_{\mathbf{Y}}\lambda_{t}^{*})^{\top}.\end{split}$ (5)
Here, $\Sigma(t,\mathbf{Y}_{t}^{*})=\bm{\sigma}(t,\mathbf{Y}_{t}^{*})\Psi\bm{\sigma}(t,\mathbf{Y}_{t}^{*})^{\top}$ is the instantaneous covariance matrix of asset returns (assuming $\bm{\sigma}$ is diagonal, this matches previous definition). The first term represents the myopic demand, driven by instantaneous expected excess return and risk aversion. The second term represents the intertemporal hedging demand, driven by the desire to hedge against unfavorable changes in the state variables $\mathbf{Y}_{t}$, depending on the sensitivity of marginal utility to state variables ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$), state variable volatility ($\bm{\sigma}_{Y}$), and correlations ($\rho,\Psi,\Phi$).
In the special case $n=1,k=1$, we have $\Sigma=\sigma^{2}$, $\Psi=1$, $\rho$ is a scalar correlation, $\bm{\sigma}=\sigma$, $\bm{\sigma}_{Y}=\sigma_{Y}$. The formula (2) simplifies to:
$\pi_{t}^{*}\;=\;-\frac{1}{X_{t}^{*}(\partial_{x}\lambda_{t}^{*})\sigma^{2}}\left\{\lambda_{t}^{*}\bigl{(}\mu(t,Y_{t}^{*})-r(t,Y_{t}^{*})\bigr{)}+\sigma\rho\sigma_{Y}(\partial_{Y}\lambda_{t}^{*})\right\},$
which clearly separates the myopic and hedging components.
@@ 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
In the preceding sections, we established the theoretical underpinnings of the continuous-time multi-asset portfolio problem using Pontryagin’s Maximum Principle (PMP). This principle provides necessary optimality conditions characterized by a coupled system of forward-backward stochastic differential equations (FBSDEs) for the state and costate variables (Section 2.2). However, finding analytical solutions to these FBSDEs is often infeasible, especially in high-dimensional settings involving numerous assets ($\bm{\pi}_{t}\in\mathbb{R}^{n}$) or complex exogenous state dynamics ($\mathbf{Y}_{t}\in\mathbb{R}^{k}$). Consequently, translating the continuous-time optimality conditions into a practical solution requires effective numerical approximations.
We adopt a gradient-based approach rooted in modern deep learning techniques. The core idea is to directly parameterize the investor’s policy, both the investment strategy $\bm{\pi}_{t}$ and the consumption rate $C_{t}$, using neural networks, denoted as $\bm{\pi}_{\theta}(t,x,\mathbf{y})$ and $C_{\phi}(t,x,\mathbf{y})$, where $(\theta,\phi)$ are the network parameters. Optimization then involves iteratively refining these parameters to maximize the investor’s expected utility.
A key challenge is computing the gradient of the expected utility objective with respect to the policy parameters $(\theta,\phi)$. Furthermore, a practical policy must be effective across a relevant range of potential starting times $t_{0}$, initial wealth levels $x_{0}$, and initial state variable values $\mathbf{y}_{0}$.
Let $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ denote the conditional expected utility obtained when starting from state $(t_{0},x_{0},\mathbf{y}_{0})$ and following the policy defined by $(\theta,\phi)$:
@@
Here, the expectation $\mathbb{E}[\cdot|\cdot]$ is taken over all possible paths of the state processes $(X_{u},\mathbf{Y}_{u})$ for $u\in[t_{0},T]$, conditional on starting at $(x_{0},\mathbf{y}_{0})$ at time $t_{0}$.
To ensure the policy performs well across various starting scenarios, we define the optimization target as the extended value function, $\widetilde{J}(\theta,\phi)$. This is the expected value of the conditional utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ when the initial state $(t_{0},x_{0},\mathbf{y}_{0})$ is drawn from a chosen distribution $\eta$ over a relevant domain $\mathcal{D}\subset[0,T]\times\mathbb{R}^{+}\times\mathbb{R}^{k}$:
$\widetilde{J}(\theta,\phi)\;=\;\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\Big{[}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\Big{]}.$ (7)
Maximizing this extended objective $\widetilde{J}(\theta,\phi)$ encourages the learning of a robust policy applicable across the domain $\mathcal{D}$ covered by the initial state distribution $\eta$.
This paper develops and implements the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, extending the core ideas of Huh, (2024) to the multi-asset, state-dependent setting, with the goal of maximizing the extended value function $\widetilde{J}(\theta,\phi)$. The PG-DPO framework leverages automatic differentiation, specifically backpropagation-through-time (BPTT), applied to Monte Carlo simulations drawn from the initial state distribution $\eta$ and evolved according to the system dynamics.
A central feature of applying BPTT in this context is that it not only provides unbiased estimates of the policy gradient $\nabla\widetilde{J}$, necessary for optimizing the neural network parameters $(\theta,\phi)$, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each simulated path. These costates are crucial as they represent the sensitivities of the expected utility to state perturbations, as dictated by PMP.
This dual output of BPTT–policy gradients and costate estimates–allows for two distinct algorithmic approaches, both offering scalable alternatives to traditional methods hindered by the curse of dimensionality:
  • The Baseline PG-DPO algorithm (§3.1) utilizes the BPTT-derived policy gradient estimates ($\nabla_{\text{batch}}\approx\nabla\widetilde{J}$) directly within a stochastic gradient ascent procedure to train the neural network policy $(\bm{\pi}_{\theta},C_{\phi})$ end-to-end. In this approach, the costate estimates primarily serve to structure the gradient calculation, implicitly guiding the optimization according to PMP.
  • The Two-Stage PG-DPO algorithm (§3.2) takes a different route. It uses the baseline procedure for a short warm-up phase specifically to generate stable and accurate estimates of the costates (potentially involving Monte Carlo averaging). These refined costate estimates are then directly plugged into the analytical first-order optimality conditions derived from PMP to define the deployed control actions, bypassing the direct use of the warm-up phase neural networks.
Both variants avoid explicit FBSDE solvers and leverage the efficiency of BPTT, but they differ significantly in how they utilize the information generated during backpropagation to arrive at the final control policy. The subsequent subsections detail each approach.

### v2

@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
For clarity on the noise structure, we define the combined $(n+k)$-dimensional Brownian vector $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$. The covariance structure of its increments $d\mathbf{W}_{t}$ is given by the matrix $\bm{\Omega}$:
$\operatorname{Cov}\bigl{(}d\mathbf{W}_{t}\bigr{)}=\bm{\Omega}\,dt=\begin{pmatrix}\Psi&\rho\\ \rho^{\top}&\Phi\end{pmatrix}\,dt,$
where $\Psi=(\psi_{ij})\in\mathbb{R}^{n\times n}$ is the covariance matrix of $d\mathbf{W}^{X}_{t}$, $\Phi=(\phi_{ij})\in\mathbb{R}^{k\times k}$ is the covariance matrix of $d\mathbf{W}^{Y}_{t}$, and $\rho=(\rho_{ij})\in\mathbb{R}^{n\times k}$ is the cross-covariance matrix.
Assumption 3 (Noise covariance).
The block covariance matrix $\bm{\Omega}=\begin{pmatrix}\Psi&\rho\\[1.79997pt] \rho^{\top}&\Phi\end{pmatrix}$ is constant in time and uniformly positive definite: there exists $\lambda_{\Omega}>0$ such that $z^{\top}\bm{\Omega}z\geq\lambda_{\Omega}\lVert z\rVert^{2}$ for all $z\in\mathbb{R}^{n+k}$. (Consequently, $\Psi$ and $\Phi$ are individually positive definite and hence invertible.)
The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ to maximize the expected discounted utility functional $J$:
[…]
subject to the wealth dynamics (1) and admissibility constraints (Assumption 2). Here $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ weights the utility of terminal wealth (bequest motive). The utility function $U:(0,\infty)\to\mathbb{R}$ represents the investor’s preferences. A widely used specification is the Constant Relative Risk Aversion (CRRA) form:
$U(x)=\begin{cases}\displaystyle\frac{x^{1-\gamma}}{1-\gamma},&\gamma>0,\gamma\neq 1,\\[6.0pt] \ln(x),&\gamma=1,\end{cases}$
with $\gamma>0$ denoting the coefficient of relative risk aversion (see Pennacchi, , 2008).
Assumption 2 (iii) bounds consumption by $C_{t}\leq\xi X_{t}$ with $\xi\in(0,1)$, while Assumption 4 (b) gives a $C^{2}$, strictly concave utility $U$ satisfying the Inada limits. Because $\delta>0$ and $X_{t}>0$ a.s. (Assumption 2 (iv)), we have
$\mathbb{E}\!\bigl{[}\,e^{-\delta t}\,U(C_{t})\bigr{]}\leq e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}<\infty,\quad\forall t\in[0,T],$
and similarly $\mathbb{E}[U(X_{T})]<\infty$. Hence the functional $J(\bm{\pi},C)$ in (2) is finite for every admissible control pair, so the value function $V$ is well defined.
The value function $V(t,x,\mathbf{y})$ associated with this problem represents the maximum expected utility achievable starting from wealth $x$ and state $\mathbf{y}$ at time $t$. Formally,
[…]
where the supremum is taken over all admissible control processes on $[t,T]$, and $\mathbb{E}_{t,x,\mathbf{y}}[\cdot]=\mathbb{E}[\cdot|X_{t}=x,\mathbf{Y}_{t}=\mathbf{y}]$ denotes the conditional expectation.
Additional assumptions are needed for the well-posedness of the problem and regularity of the value function.
Assumption 4 (Uniform non-degeneracy and utility).
  1. (a)
    Uniform non-degeneracy of risky-asset volatility. Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$ be the instantaneous covariance matrix of the risky asset returns (Note: original used $\bm{\sigma}\bm{\sigma}^{\top}$, assuming $\Psi=I$; using $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ is more general if $\mathbf{W}^{X}$ components are correlated and $\bm{\sigma}$ is diagonal). We assume there exists $\lambda>0$ such that
    $\mathbf{z}^{\top}a(t,\mathbf{y})\,\mathbf{z}\;\geq\;\lambda\,\lVert\mathbf{z}\rVert^{2},\quad\forall(t,\mathbf{y},\mathbf{z})\in[0,T]\times\mathbb{R}^{k}\times\mathbb{R}^{n}.$
    (Equivalently, the covariance matrix of risky returns $a(t,\mathbf{y})$ is uniformly positive definite, ensuring invertibility, which is crucial for solving the portfolio problem.)
  2. (b)
    Utility specification. The utility function $U:(0,\infty)\to\mathbb{R}$ satisfies $U\in C^{2}(0,\infty)$, $U^{\prime}>0,\;U^{\prime\prime}<0$ (strictly increasing and concave), and the Inada conditions $\lim_{c\downarrow 0}U^{\prime}(c)=\infty,\;\lim_{c\uparrow\infty}U^{\prime}(c)=0.$
  3. (c)
    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$ and the terminal weight satisfies $\kappa\geq 0$.
Assumption 4 (a) guarantees the diffusion related to risky assets is non-degenerate, a key ingredient for the regularity of the associated Hamilton-Jacobi-Bellman (HJB) PDE. Parts (b)-(c) ensure the objective functional is well-defined and satisfies standard preference assumptions. Combined with the coefficient regularity in Assumption 1, these conditions typically yield the existence, uniqueness, and sufficient smoothness (e.g., $V\in C^{1,2,2}$) of the solution to the HJB equation, allowing for verification theorems to be applied; see Oksendal, (2013, Chs. 6-7) and related literature on optimal portfolio choice.
In our numerical approach, we parameterize the control policies using neural networks (NNs). The portfolio allocation $\bm{\pi}_{t}$ and consumption rate $C_{t}$ are represented as functions of the current time $t$, wealth $X_{t}$, and state $\mathbf{Y}_{t}$, determined by the NN parameters $\theta$ and $\phi$:
$\bm{\pi}_{t}=\bm{\pi}_{\theta}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)},\quad C_{t}=C_{\phi}\!\bigl{(}t,\,X_{t},\,\mathbf{Y}_{t}\bigr{)}.$
These parameters $\theta$ and $\phi$ are then optimized to maximize the objective $J$.
2.2 Optimality Conditions via Pontryagin’s Maximum Principle
We now apply Pontryagin’s Maximum Principle (PMP) to the continuous-time portfolio problem formulated in Section 2.1, where market parameters depend on the exogenous state process $\{\mathbf{Y}_{t}\}_{0\leq t\leq T}$. PMP provides necessary conditions for optimality through a system of forward-backward stochastic differential equations (FBSDEs), which couple the dynamics of the state variables with those of associated costate variables.
The system state includes the investor’s wealth $X_{t}$ and the exogenous state variables $\mathbf{Y}_{t}\in\mathbb{R}^{k}$. The corresponding costate processes represent the sensitivity of the value function $V$ to changes in the state: $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ is the scalar costate for wealth (marginal value of wealth), and $\bm{\eta}_{t}=V_{\mathbf{y}}(t,X_{t},\mathbf{Y}_{t})$ is the $k$-dimensional costate vector for the exogenous state variables (marginal value of state variables).
Under an optimal control policy $\bigl{(}\bm{\pi}_{t}^{*},\,C_{t}^{*}\bigr{)}$, the evolution of the system state, comprising the wealth process $X_{t}^{*}$ and the exogenous variables $\mathbf{Y}_{t}$, is described by the following forward stochastic differential equations:
$dX_{t}^{*}$ $=\Bigl{[}X_{t}^{*}\,\Bigl{(}r\bigl{(}t,\mathbf{Y}_{t}\bigr{)}+\bm{\pi}_{t}^{*\top}\,\bigl{(}\bm{\mu}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}\Bigr{)}-C_{t}^{*}\Bigr{]}\,dt$
$\quad+X_{t}^{*}\,\bm{\pi}_{t}^{*\top}\,\bm{\sigma}(t,\mathbf{Y}_{t})\,d\mathbf{W}_{t}^{X},$ (3)
$d\mathbf{Y}_{t}$ $=\bm{\mu}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,dt+\bm{\sigma}_{Y}\bigl{(}t,\mathbf{Y}_{t}\bigr{)}\,d\mathbf{W}_{t}^{Y},$ (4)
with initial conditions $X_{0}^{*}=x_{0}>0$ and $\mathbf{Y}_{0}=\mathbf{y}_{0}\in\mathbb{R}^{k}$. Recall the definitions of the volatility matrix $\bm{\sigma}(t,\mathbf{Y}_{t})$ and the covariance structure $\bm{\Omega}$ of the combined Brownian motion $\mathbf{W}_{t}=(\mathbf{W}^{X\top}_{t},\mathbf{W}^{Y\top}_{t})^{\top}$.
Complementing these forward dynamics, the optimal costate processes $(\lambda_{t}^{*},\bm{\eta}_{t}^{*})$ satisfy a system of backward stochastic differential equations (BSDEs). These equations are driven by derivatives of the Hamiltonian $\mathcal{H}$ associated with the control problem:
$-d\lambda_{t}^{*}$ $=\mathcal{H}_{x}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\lambda X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\lambda Y}\,d\mathbf{W}_{t}^{Y},$ (5)
$-d\bm{\eta}_{t}^{*}$ $=\mathcal{H}_{\mathbf{y}}\bigl{(}t,X_{t}^{*},\mathbf{Y}_{t},\bm{\pi}_{t}^{*},C_{t}^{*},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}\bigr{)}\,dt-\mathbf{Z}_{t}^{\eta X}\,d\mathbf{W}_{t}^{X}-\mathbf{Z}_{t}^{\eta Y}\,d\mathbf{W}_{t}^{Y}.$ (6)
Here, $\mathcal{H}_{x}=\partial\mathcal{H}/\partial x$ and $\mathcal{H}_{\mathbf{y}}=\nabla_{\mathbf{y}}\mathcal{H}$ represent the partial derivatives of the Hamiltonian. The terms $\mathbf{Z}_{t}^{\lambda X}\in\mathbb{R}^{1\times n}$, $\mathbf{Z}_{t}^{\lambda Y}\in\mathbb{R}^{1\times k}$, $\mathbf{Z}_{t}^{\eta X}\in\mathbb{R}^{k\times n}$, and $\mathbf{Z}_{t}^{\eta Y}\in\mathbb{R}^{k\times k}$ are adapted processes arising from the martingale representation theorem applied to the costates, collectively denoted by $\mathbf{Z}_{t}$. The derivatives $\mathcal{H}_{x},\mathcal{H}_{\mathbf{y}}$ generally depend on the state, costate, controls, and the $\mathbf{Z}_{t}$ terms themselves, reflecting the system’s interconnectedness through the value function’s second-order derivatives. The terminal conditions are determined by the objective functional: $\lambda_{T}^{*}=\kappa\,e^{-\delta T}\,U^{\prime}\bigl{(}X_{T}^{*}\bigr{)}$ and $\bm{\eta}_{T}^{*}=\mathbf{0}$.
The application of PMP involves maximizing a Hamiltonian function. To properly account for the dependence of the wealth diffusion on the control $\bm{\pi}$, the Hamiltonian should incorporate the interaction between the costate dynamics (represented by $\mathbf{Z}$ terms) and the diffusion coefficients. Following formulations in stochastic control theory (e.g., Yong & Zhou, , 2012), the appropriate Hamiltonian $\mathcal{H}$ can be expressed as:
@@ 2.2 Optimality Conditions via Pontryagin’s Maximum Principle
$\begin{split}&\mathcal{H}\bigl{(}t,x,\mathbf{y},\bm{\pi},C,\lambda,\bm{\eta},\mathbf{Z}^{\lambda X},\mathbf{Z}^{\lambda Y},\mathbf{Z}^{\eta X},\mathbf{Z}^{\eta Y}\bigr{)}\\ &\quad=e^{-\delta t}\,U(C)+\lambda\Bigl{[}x\bigl{(}r(t,\mathbf{y})+\bm{\pi}^{\top}(\bm{\mu}(t,\mathbf{y})-r(t,\mathbf{y})\mathbf{1})\bigr{)}-C\Bigr{]}+\bm{\eta}^{\top}\,\bm{\mu}_{Y}(t,\mathbf{y})\\ &\quad\quad+\mathbf{Z}^{\lambda X}\bigl{(}x\,\bm{\sigma}(t,\mathbf{y})^{\top}\bm{\pi}\bigr{)}+\text{Tr}\bigl{(}\mathbf{Z}^{\eta Y\top}\bm{\sigma}_{Y}(t,\mathbf{y})\bigr{)}+\dots\end{split}$ (7)
For clarity in deriving the FOC for $\bm{\pi}$, the crucial term involving $\bm{\pi}$ beyond the drift component is the one coupling $\mathbf{Z}^{\lambda X}$ with the wealth diffusion $x\,\bm{\pi}^{\top}\bm{\sigma}(t,\mathbf{Y}_{t})$ originating from $d\mathbf{W}_{t}^{X}$. We focus on these terms for the maximization step.
To ensure the well-posedness of this FBSDE system and the applicability of PMP, additional smoothness conditions on the model coefficients are typically required.
Assumption 5 (Additional regularity for PMP).
In addition to Assumptions 1, 2, and 4, we assume:
  1. (a)
    $C^{1}$-regular coefficients w.r.t. $\mathbf{y}$. All coefficient functions $\bm{\mu}_{Y},\,\bm{\sigma}_{Y},\,\bm{\mu},\,\bm{\sigma}\,$ $r$ are continuously differentiable with respect to the state vector $\mathbf{y}$. Furthermore, their partial derivatives ($\partial_{\mathbf{y}}f$) are globally Lipschitz continuous in $\mathbf{y}$ and satisfy linear growth conditions. This ensures that the drift terms in the costate BSDEs, involving derivatives like $\mathcal{H}_{\mathbf{y}}$, satisfy the required regularity conditions for the existence and uniqueness of solutions (cf. Yong & Zhou, , 2012).
The core requirement of Pontryagin’s Maximum Principle is that the optimal controls $(\bm{\pi}_{t}^{*},C_{t}^{*})$ must maximize the Hamiltonian $\mathcal{H}$ (7) pointwise in time, almost surely, given the optimal state and costate processes $(X_{t}^{*},\mathbf{Y}_{t},\lambda_{t}^{*},\bm{\eta}_{t}^{*},\mathbf{Z}_{t}^{*})$. Assuming interior solutions allows us to derive the first-order conditions (FOCs) for optimality. Maximization with respect to the consumption rate $C_{t}$ yields the standard condition:
$e^{-\delta t}\,U^{\prime}(C_{t}^{*})=\lambda_{t}^{*}.$ (8)
This classic result links optimal consumption $C_{t}^{*}=(U^{\prime})^{-1}(e^{\delta t}\lambda_{t}^{*})$ inversely to the marginal value of wealth $\lambda_{t}^{*}$. Maximization with respect to the portfolio weights $\bm{\pi}_{t}$ requires differentiating $\mathcal{H}$ with respect to $\bm{\pi}$, including the term involving $\mathbf{Z}^{\lambda X}$. From (7), the relevant terms are $\lambda x\bm{\pi}^{\top}(\bm{\mu}-r\mathbf{1})$ and $\mathbf{Z}^{\lambda X}(x\,\bm{\sigma}^{\top}\bm{\pi})$. Setting the gradient $\nabla_{\bm{\pi}}\mathcal{H}=\mathbf{0}$ gives (assuming $X_{t}^{*}>0$):
$\lambda_{t}^{*}X_{t}^{*}(\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1})+X_{t}^{*}\bm{\sigma}(t,\mathbf{Y}_{t})(\mathbf{Z}_{t}^{\lambda X*})^{\top}=\mathbf{0}.$
Dividing by $X_{t}^{*}$, we obtain the FOC for the optimal portfolio:
[…]
This first-order condition implicitly defines the optimal portfolio $\bm{\pi}_{t}^{*}$ through its dependence on the costate $\lambda_{t}^{*}$ and the martingale component $\mathbf{Z}_{t}^{\lambda X*}$.
To obtain an explicit expression for $\bm{\pi}_{t}^{*}$, we need to relate the martingale component $\mathbf{Z}^{\lambda X*}$ to the derivatives of the value function. Applying Ito’s lemma to $\lambda_{t}=V_{x}(t,X_{t},\mathbf{Y}_{t})$ and matching terms with the BSDE (5) establishes this relationship (the precise derivation often involves careful application of Ito’s formula for processes driven by correlated Brownian motions, see e.g., standard texts or appendices like Appendix A). This procedure yields an expression for $\mathbf{Z}^{\lambda X*}$ in terms of $V_{xx}$, $V_{x\mathbf{y}}$, the optimal controls $\bm{\pi}_{t}^{*}$, and the model parameters ($\bm{\sigma},\bm{\sigma}_{Y},\Psi,\rho$). Substituting this expression into the FOC (9) and solving the resulting algebraic equation for $\bm{\pi}_{t}^{*}$ leads to the well-known optimal portfolio formula:
$\bm{\pi}_{t}^{*}$ $=-\frac{V_{x}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bigl{(}\bm{\mu}(t,\mathbf{Y}_{t})-r(t,\mathbf{Y}_{t})\mathbf{1}\bigr{)}$ (10)
$\qquad\qquad-\frac{V_{x\mathbf{y}}}{X_{t}V_{xx}}\,\Sigma^{-1}(t,\mathbf{Y}_{t})\,\bm{\sigma}(t,\mathbf{Y}_{t})\,\rho\,\bm{\sigma}_{Y}^{\!\top}(t,\mathbf{Y}_{t}).$
Here, $\Sigma(t,\mathbf{Y}_{t})=\bm{\sigma}(t,\mathbf{Y}_{t})\Psi\bm{\sigma}(t,\mathbf{Y}_{t})^{\top}$ represents the instantaneous covariance matrix of the risky asset returns. The value function derivatives $V_{x}=\lambda_{t}^{*}$, $V_{xx}=\partial_{x}\lambda_{t}^{*}$, and $V_{x\mathbf{y}}=\partial_{\mathbf{y}}\lambda_{t}^{*}$ are evaluated along the optimal path $(t,X_{t}^{*},\mathbf{Y}_{t})$. The formula clearly separates the portfolio into two components: the myopic demand, which optimizes the immediate risk-return trade-off as if the investment opportunity set were constant, and the intertemporal hedging demand, which adjusts the portfolio to mitigate risks arising from future stochastic changes in the state variables $\mathbf{Y}_{t}$. The hedging demand’s magnitude and direction depend on the correlation $\rho$ between asset returns and state variable innovations, and crucially on $V_{x\mathbf{y}}$, which measures how the marginal value of wealth changes with the state variables.
For illustration, in the simplified case with one risky asset ($n=1$) and one state variable ($k=1$), the optimal allocation (10) becomes:
$\pi_{t}^{*}=-\frac{V_{x}}{X_{t}V_{xx}\sigma^{2}}\bigl{(}\mu(t,Y_{t})-r(t,Y_{t})\bigr{)}-\frac{V_{xy}}{X_{t}V_{xx}\sigma^{2}}\sigma\rho\sigma_{Y},$
clearly showing the separation between the myopic component, proportional to the market price of risk, and the hedging component, proportional to $V_{xy}$ and the correlation $\rho$.
@@ 3 A Gradient-Based Algorithm for Multi-Asset Policy Optimization
While Pontryagin’s Maximum Principle (PMP) provides necessary optimality conditions via forward-backward stochastic differential equations (FBSDEs, Section 2.2), solving these analytically is often infeasible in high dimensions. We therefore adopt a gradient-based numerical approach using deep learning. The core idea is to parameterize the investor’s policy $(\bm{\pi}_{t},C_{t})$ with neural networks $(\bm{\pi}_{\theta}(t,x,\mathbf{y}),C_{\phi}(t,x,\mathbf{y}))$, where $(\theta,\phi)$ are the parameters to be optimized.
The objective is to find parameters $(\theta,\phi)$ that maximize the expected utility achieved by the policy $(\bm{\pi}_{\theta},C_{\phi})$. To ensure robustness across different starting scenarios, we aim to maximize the extended value function, $\widetilde{J}(\theta,\phi)$. This is defined as the expected value of the conditional expected utility $J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)$ achieved under policy $(\theta,\phi)$, where the initial state $(t_{0},x_{0},\mathbf{y}_{0})$ is drawn from a distribution $\eta$ over a relevant domain $\mathcal{D}$:
@@
$\widetilde{J}(\theta,\phi)=\mathbb{E}_{(t_{0},x_{0},\mathbf{y}_{0})\,\sim\,\eta}\Big{[}J(t_{0},x_{0},\mathbf{y}_{0};\theta,\phi)\Big{]}$ (12)
Here, $\mathbb{E}^{\theta,\phi}[\cdot|\cdot]$ denotes expectation under the dynamics generated by policy $(\theta,\phi)$.
This paper develops and implements the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework to maximize $\widetilde{J}(\theta,\phi)$. PG-DPO leverages automatic differentiation, specifically backpropagation-through-time (BPTT), applied to Monte Carlo simulations drawn from the initial state distribution $\eta$. A key feature is that BPTT applied to these simulations yields not only unbiased estimates of the policy gradient $\nabla\widetilde{J}$ needed for optimization, but also inherently calculates estimates of the Pontryagin costate variables ($\lambda_{k}^{(i)},\bm{\eta}_{k}^{(i)}$) along each path, representing sensitivities dictated by PMP.
This dual output enables distinct algorithmic approaches detailed in the subsequent subsections.